On a cold night the temperatures of the two surfaces of a glass windowpane in a house are 7.8° C on the outer surface and 8.0° C on the inner surface when the air temperature is 0° C outside and 21.2° C inside. The glass is 3.33 mm thick. (The thermal conductivity for glass is 0.8 W/m-C°.)

(a) Find the rate at which heat is conducted through the glass per square meter of window area.


(b) Find the value of the convection coefficients for the heating of the inner surface of the glass and for the cooling of the outer surface.

houter = _____ W/m2-C°
hinner = _____ W/m2-C°
Answer #1

a) rate at which the heat is transferred = kA dT/dx

Where k = thermal conductivity of the material

A = cross sectional area

dt/dx = temperature gradient

Q = 0.8 * 1 * (8-7.8) / (3.33 * 10^-3)

Q = 48.048 W

B) this heat transferred between the plates = convective heat transferred from the outer surface to the glass

48.048 = hA dt

Where h = convective heat transfer coefficient

dt = change in temperature

48.048 = h*1*(7.8-0)

h = 6.16 W/m2 C

Convective heat transfer coefficient for outer surface

48.048 = h*1 * (21.2 -8)

h = 3.64 W/m2C